Circular No. 7717

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V445 PUPPIS M. P. Rupen, V. Dhawan, and A. Mioduszewski, National Radio As-

tronomy Observatory, report: “Observations with the Very Large Array (VLA) at 8.46 GHz on Sept. 9 show a 9.6-mJy point source (rms 0.095 mJy/beam) at α = 7^{h}37^{m}56^{s}.87, δ = −25^{o}56′58′′.9 (equinox 2000.0; uncertainty 0′′.1–0′′.2 in each coordinate). This agrees with the optical position of V445 Pup (IAUC 7556). Observations at the same frequency on Jan. 18 gave no detection (rms 0.095 mJy/beam). VLA data taken on Sept. 11 and 12 confirm the detection and show a slowly decaying source (exponential timescale of 25 ± 3 days) with a power-law spectrum (flux density ∝ ν^{−0.7}, where ν is the frequency) between 5 and 15 GHz. The 1.425-GHz flux density, however, is lower than a simple extrapolation would suggest, giving an inferred foreground opacity of 0.4 at that frequency. If this is correct, either the source is small enough (milliarcsec) for synchrotron self-absorption, or there is a great deal of ionized gas providing free-free absorption along the line-of-sight. The radio emission probably comes either from a recent energetic event, or from the development of a shock associated with the Dec. 2000 outburst. Observations at optical, infrared, and x-ray wavelengths would be very useful in tracing the origin of the radio emission.”

R. M. Wagner, LBT Observatory; G. Schwarz, Steward Observatory; S. G. Starrfield, Arizona State University; and C. B. Foltz, MMT Observatory, write: “Spectra (range 400–950 nm, resolution 0.85 nm) of V445 Pup, obtained on Apr. 15–16 UT (just prior to conjunction with the sun) using the Bok 2.3-m telescope on Kitt Peak, show strengthening emission lines of carbon compared to our previous reports (cf. IAUC 7556, 7571). The emission lines include C I (multiplets 3, 9–10) and C II (multiplets 1–3, 5–6, 14–16, 19–20), as well as features arising from Fe II, O I, Mg I, Mg II, N I, Ca I, Ca II, and perhaps Na I (described previously). Emission lines arising from He I (587.5, 667.8, 706.5, and 728.1 nm) are also present (cf. IAUC 7556). The equivalent width of the C II line at 658 nm was 4.6 nm. A spectrum (range 390–540 nm, resolution 0.14 nm) obtained on Mar. 29 by S. Howell and P. Szkody with the 6.5-m MMT shows that many of these features exhibit P-Cyg profiles in which the absorption components are relatively sharp (FWHM 130 km/s; terminal velocity −600 km/s), while the emission components exhibit considerable structure and long red tails extending to ∼ +3800 km/s, suggesting an asymmetric outburst. A broad and weak emission feature (equivalent width 0.32 nm) at 486.1 nm might be Hβ. Additional spectroscopic observations are urgently requested.”

2001 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green